Cindy had $6.05 in quarters and nickels. How many of each coin does she have if there are 53 coins in all?
Answer by stanbon(75887)

Let number of nickels be "x".
Value of nickels is "5x" cents
Number of quarters is "53-x"
Value of quarters is 25(53-x)="1325-25x" cents
EQUATION:
value + value = 605 cents
5x+1325-25x=605
-20x=-720
x=36 (number of nickels)
53-x=17 (number of quarters)
Cheers,
stan H.